### v1

@@ Abstract
Recent advancements in large-scale video-language models have shown significant potential for real-time planning and detailed interactions. However, their high computational demands and the scarcity of annotated datasets limit their practicality for academic researchers. In this work, we introduce VideoLLaMB, a novel framework that utilizes temporal memory tokens within bridge layers to allow for the encoding of entire video sequences alongside historical visual data, effectively preserving semantic continuity and enhancing model performance across various tasks. This approach includes recurrent memory tokens and a SceneTilling algorithm, which segments videos into independent semantic units to preserve semantic integrity. Empirically, VideoLLaMB significantly outstrips existing video-language models, demonstrating a $5.5$ points improvement over its competitors across three VideoQA benchmarks, and $2.06$ points on egocentric planning. Comprehensive results on the MVBench show that VideoLLaMB-7B achieves markedly better results than previous 7B models of same LLM. Remarkably, it maintains robust performance as PLLaVA even as video length increases up to 8×8$\times$8 ×. Besides, the frame retrieval results on our specialized Needle in a Video Haystack (NIAVH) benchmark, further validate VideoLLaMB’s prowess in accurately identifying specific frames within lengthy videos. Our SceneTilling algorithm also enables the generation of streaming video captions directly, without necessitating additional training. In terms of efficiency, VideoLLaMB, trained on 16 frames, supports up to 320 frames on a single Nvidia A100 GPU with linear GPU memory scaling, ensuring both high performance and cost-effectiveness, thereby setting a new foundation for long-form video-language models in both academic and practical applications.
[…]
@@ 1 Introduction
The recent advances of large-scale video language models, represented with GPT4-o [^1] https://openai.com/index/hello-gpt-4o/ and Project Astra [^2] https://deepmind.google/technologies/gemini/project-astra/, have amazed the world by their potential for nuanced interaction with the real world environment, particularly for real-time planning that demands the ability to observe the current state and draw from long-term memory. However, training such super-scale video-language foundational models is infeasible to academic researchers because of the massive computational cost required by the complex, high-dimensional nature of long video data, coupled with a scarcity of well-annotated, public video-language datasets, which poses significant challenges for extensively scaling video-language models akin to those observed with large language models (LLMs).
@@
To circumvent these challenges, the community has witnessed a growing interest in developing computationally efficient multimodal large language models (MLLMs). Traditional methods resort to video compression strategies, such as sampling (Zhang et al., 2023b; Lin et al., 2023), aggregation (Xu et al., 2024b), semantic consolidation (Song et al., 2023b), and resampling (Ma et al., 2023; He et al., 2024), in order to temporally reduce the length of the video. Yet, these methods often lead to the loss of critical visual cues, undermining the model’s ability to capture essential cues. Some other approaches employ a sliding window mechanism (Qian et al., 2024b), segmenting videos into shorter clips to mitigate the computational load of processing long videos. However, segmentation can disrupt the semantic flow of content, complicating the encoding process and potentially impacting the general understanding of the video narrative. Lastly, prevalent video understanding benchmarks, primarily based on linguistic question-answering pairs, exhibit static (Lei et al., 2023) and/or language biases (Ruggeri & Nozza, 2023; Zhou et al., 2022). These biases favor models that rely more on static imagery or textual elements, respectively, and fail to provide a comprehensive assessment of a model’s capability on extended video sequences. Refer to §$\S$4 for detailed discussions.
@@
Figure 1: Overview of VideoLLaMB. We first extract the video features using an off-the-shelf vision encoder, then apply SceneTilling to segment the video into semantic segments (§$\S$2.1). Next, we use recurrent memory on these semantic segments to store video information within memory tokens (§$\S$2.2). We further employ a retrieval mechanism to update the memory tokens and address long-dependency issues (§$\S$2.3). Finally, we project the memory-token-augmented features from the current video segment into the LLM.
@@
To address these multifaceted limitations, we introduce VideoLLaMB, an innovative framework that learns temporal Memory tokens within Bridge layers that recursively encode the entire video content, ensuring that no visual information is discarded deliberately (Figure 1; §$\S$2). Specifically, we devise Memory Bridge Layers, equipped with recurrent memory tokens, function without altering the architecture of the visual encoder and LLM. Furthermore, to mitigate the risk of vanishing gradients, we maintain long-term dependencies by preserving recurrent memory tokens in a memory cache, which is periodically refreshed through a retrieval process. To compensate for the limitations of the sliding window technique, we propose SceneTilling algorithm that divides the video into relatively independent sequences of semantic segments. This reduces the dimensions within each semantic unit without sacrificing semantic details. By constructing our recurrent memory with a retrieval mechanism based on these semantic segments, our method strikes a balance between effective and efficient comprehension of the current state and long-term memory retention.
@@
In §$\S$3, we highlight the empirical advantages of VideoLLaMB in comparison with prior arts as:
  • Comprehensive long video understanding. We demonstrate the effectiveness of VideoLLaMB using two long-form video QA benchmarks: EgoSchema (Mangalam et al., 2023) and NexTQA (Xiao et al., 2021). Our results show an average improvement of $5.5$ accuracy points over PLLaVA (Xu et al., 2024b), a model with the same initialization and training video dataset. Furthermore, VideoLLaMB maintains its performance even when the video length extends to $8$ times longer than the original. Additionally, performance on MVBench (Li et al., 2023b) indicates that VideoLLaMB significantly outperforms prior models like PLLaVA using the same training data and LLM baseline.
  • Training-free streaming captioning. By employing the SceneTilling algorithm, our method can automatically predict the end of a caption in streaming video without relying on special tokens during the training phase.
@@
  • Memory-based egocentric planning. To evaluate our model’s performance in video planning tasks, we used the planning dataset EgoPlan (Chen et al., 2023). Our method achieves the best performance among all 7B video-language models, showing an improvement of $2.06$ accuracy points over PLLaVA.
  • Enhanced frame retrieval in long videos. To evaluate our model’s ability in frame retrieval for long videos, we propose a multimodal Needle in a Video Haystack (NIAVH) test. This test requires the model to predict the true answer about an inserted image in a long video. In our NIAVH pressure test, which ranges from $1$ to $320$ seconds in length, VideoLLaMB consistently retrieves the correct image needles at various depths, outperforming other methods as video length increases.
@@ 2 VideoLLaMB
VideoLLaMB is an extensible framework designed to enhance long video understanding, composed of three key modules: semantic-based segmenter (§$\S$2.1), recurrent memory layer (§$\S$2.2), and memory retriever (§$\S$2.3). Each of these components will be detailed in the subsequent sections. Figure 1 depicts the overall framework.
@@ 2.1 SceneTiling: Segmentation with Semantics
Semantic segmentation along temporal sequence has long been recognized as an important task because it preserves the non-linear structure of context and greatly aids in compressing extensive context (Rao et al., 2020; Chen et al., 2021; Mun et al., 2022; Huang et al., 2020; Wang et al., 2023d). To address the disruption of semantic flow (see §$\S$1), we introduce SceneTilling, a model-free scene segmentation algorithm inspired by TextTiling (Hearst, 1997). SceneTilling divides the entire video sequence into segments that are semantically distinct, ensuring inter-segment coherence.
@@
Aside from temporal semantic segmentation, SceneTiling enables streaming video captioning without requiring training with special tokens (Chen et al., 2024b; Zhou et al., 2024; Fu et al., 2024) (Figure 4).
@@ 2.2 Recurrent Memory Bridge Layers
Traditional recurrent memory-based Transformers (Bulatov et al., 2022; 2023; Kuratov et al., 2024) incur significant computational costs when scaled up, i.e., $\mathcal{O}(LK)$, where $L$ is the context length and $K$ is the number of segment, primarily due to its recurrent mechanism over the whole language model. More recently, some works empirically identify that linear projection best withstands visual information within MLLMs (Liu et al., 2023b; a; Zhang et al., 2024b), albeit with high space complexity, whereas the resampler has strong compressing ability on semantic information (Li et al., 2022), though it tends to miss detailed information (Xu et al., 2024a).
In this work, we devised a novel Recurrent Memory Bridge Layer, implemented as a multi-layer Transformer block, that integrates recurrent memory tokens within bridge layers to enhance the linear layer’s memorization ability. Formally, for each video segment $s_{i}$, we prepend a fixed number of memory tokens, denoted as $[m_{i};s_{i}]$, where $m_{i}$ represents the memory tokens. Subsequently, we apply standard self-attention to this sequence, yielding $[m_{i+1};o_{i}]={\rm BridgeLayer}([m_{i};s_{i}])$. Here, $m_{i+1}$ is the updated memory token, and $o_{i}$ is the visual representation from the bridge layers. This process is carried out recursively, traversing the semantic video segments while updating the memory tokens. After a total of $k$ steps, this output represents the condensed visual summary of the video sequence and will be used as the input for the LLM. As such, the Memory Bridge can compress past video into memory tokens while preserving current video scenes through projection without losing detailed information by compressing.
@@ 2.3 Memory Cache with Retrieval
One of the primary challenges associated with recurrent memory bridge layers is the potential for gradient vanishing, which can impede the model’s ability to learn long-range dependencies. To mitigate this issue, we propose the incorporation of a memory cache with a retrieval strategy designed to preserve previous states of memory.
@@
At each timestep $i$, the system stores all previous memory tokens in a memory cache, denoted as $M_{i}=[m_{1},\ldots,m_{i}]$. We employ a self-retrieval mechanism to update the current memory token $m_{i}$. Specifically, we treat $m_{i}$ as a query and the concatenated memory cache $M_{i}$ as key and value. The model performs a standard multi-head cross-attention operation to integrate information from previous timesteps into the current memory state, yielding the updated memory token
$m_{i+1}=\text{Softmax}\left(\frac{W_{i}^{Q}m_{i}(W_{i}^{K}M_{i})^{\top}}{\sqrt{d_{k}}}\right)W_{i}^{V}M_{i},$ (1)
where $W_{i}^{Q},W_{i}^{K},W_{i}^{V}$ are weight martices for query, key and value, repsectively.
@@ 3.1 Setup
We utilize Vicuna-7B-v1.5333https://huggingface.co/lmsys/vicuna-7b-v1.5 as the LLM and ViT-L/14 as the visual backbone following VideoLLava (Lin et al., 2023). Each frame is resized and cropped to a dimension of 224×224. The Memory Bridge Layers are based on a single-layer Transformer. Our model is trained with 16 frames and 4 segments, following the same video data protocol as PLLaVA. For the NIVAH test, we use the memory tokens as input to LLMs to evaluate their memory capabilities. For additional implementation details, please refer to Appendix B.
@@ 3.2 Long-form Video Understanding
We compare two model types: retrieval-based methods and generative video-language models, as discussed in Section 4. For fairness, we primarily compare against SOTA models LLaVA-NeXT-Video-DPO (Zhang et al., 2024b) and PLLaVA (Xu et al., 2024b), which use the same base model and video datasets as ours. Other works like MovieChat (Song et al., 2023b) and MA-LMM (He et al., 2024) are excluded due to inconsistent model configurations and benchmark settings (see Appendix B.3).
@@
VideoLLaVA 2023 Vicuna-7B 8 40.2
LLaVA-NeXT-Video-DPO 2024b Vicuna-7B 32 41.6
@@
Table 1: Results on Subset of EgoSchema under zero-shot setting. indicates that the model has been fine-tuned using the training data from EgoSchema.
@@
LLaMA-VID 2023d 53.8 60.0 73.0 59.5
VideoLLaVA 2023 56.9 61.0 75.0 61.3
LLaVA-NeXT-Video-DPO 2024b 55.6 61.0 73.9 61.3
@@
Method Vision Encoder LLM Size AS AP AA FA UA OE OI OS MD AL ST AC MC MA SC FP CO EN ER CI Avg.
@@
LLaVA 2023b ViT-L 7B 28.0 39.5 63.0 30.5 39.0 53.0 41.0 41.5 23.0 20.5 45.0 34.0 20.5 38.5 47.0 25.0 36.0 27.0 26.5 42.0 36.0
Video-LLaMA 2023b CLIP-G 7B 27.5 25.5 51.0 29.0 39.0 48.0 40.5 38.0 22.5 22.5 43.0 34.0 22.5 32.5 45.5 32.5 40.0 30.0 21.0 37.0 34.1
LLaMA-Adapter 2023d ViT-B 7B 23.0 28.0 51.0 30.0 33.0 53.5 32.5 33.5 25.5 21.5 30.5 29.0 22.5 41.5 39.5 25.0 31.5 22.5 28.0 32.0 31.7
Video-ChatGPT 2023 ViT-L 7B 23.5 26.0 62.0 22.5 26.5 54.0 28.0 40.0 23.0 20.0 31.0 30.5 25.5 39.5 48.5 29.0 33.0 29.5 26.0 35.5 32.7
VideoChat 2023c CLIP-G 7B 33.5 26.5 56.0 33.5 40.5 53.0 40.5 30.0 25.5 27.0 48.5 35.0 20.5 42.5 46.0 26.5 41.0 23.5 23.5 36.0 35.5
VideoChat2β 2023b UMT-L 7B 66.0 47.5 83.5 49.5 60.0 58.0 71.5 42.5 23.0 23.0 88.5 39.0 42.0 58.5 44.0 49.0 36.5 35.0 40.5 65.5 51.1
@@
Table 3: Results on MVBench (Li et al., 2023b) multi-choice question answering. We list GPT-4V in the first row group as a reference. The second row group includes image-based MLLMs. The third row group includes video-based MLLMs. We highlight top-3 results among all 7B models of each category in purple. $\alpha$: training data from Xu et al. (2024b). $\beta$: training with data from Li et al. (2023b).
@@
EgoSchema (Mangalam et al., 2023) consists of egocentric videos, each averaging 180 seconds in length. This video QA dataset focuses on aspects such as understanding, reasoning, and long-term memory. In our experiment, we follow the precedent set by previous studies and use the public subset for evaluation. The results are presented in Table 2. Overall, our method significantly outperforms current generative video language models trained on similar data, demonstrating robust performance compared to other approaches and confirming its efficacy. Specifically, we compare our method with PLLaVA Xu et al. (2024b), which shares the same training data, LLM backbones, and input number of frames. Our method shows significant improvements over PLLaVA, indicating its superiority in understanding long egocentric videos. While our method does not yet match the performance of fine-tuned retrieval-based methods, we plan to apply our approach to larger language models to bridge this performance gap.
@@
The model is trained on 16-frame sequences, divided into 4 segments. However, in real-world scenarios, videos can be significantly longer than this training configuration. To demonstrate VideoLLaMB’s ability to extrapolate to longer videos, we conducted experiments on EgoSchema under two conditions: 1) dynamic segments, which adaptively control the number of segments based on the SceneTilling threshold, and 2) static segments, fixed at 4 segments. Results in Figure 2 reveal that dynamic segments are more effective than static segments, especially for shorter videos, indicating that our method can effectively maintain an appropriate number of segments. However, as video length increases, the performance of dynamic segments declines, notably at the 32-frame mark, where both strategies use four segments. Beyond this point, increasing the number of segments results in diminishing returns, likely due to the domain gap from training on shorter videos. To address this issue, we plan to fine-tune our models on longer videos for more substantial improvements. Overall, compared to PLLaVA, our method maintains consistent performance as the input length increases. In summary, our approach effectively extracts key information from videos, outperforming the simple pooling strategies used for memory consolidation in existing methods.
Results on NExTQA (Xiao et al., 2021)
@@
NExTQA (Xiao et al., 2021), featuring daily-life videos that average 45 seconds in length, is designed to test a variety of question types, specifically temporal, causal, and descriptive questions. We applied our method to NExTQA to evaluate its temporal grounding ability. To maintain consistency with established benchmarks, we used the validation set for evaluation. In Table 2, we present the comprehensive results of our analysis. For a fair comparison, our primary benchmark is against PLLaVA, which includes instruction data from the NExTQA training set. Our method surpasses PLLaVA by $2.9$ points. Notably, our approach demonstrates a significant enhancement in the temporal setting, achieving a $4.6$ point improvement over PLLaVA. These results indicate that our scene-segment aware method effectively improves the model’s temporal grounding ability by compressing abundant information within scenes that share high semantic similarity.
Results on Comprehensive Video Understanding Benchmark

### v2

@@ Abstract
Recent advancements in large-scale video-language models have shown significant potential for real-time planning and detailed interactions. However, their high computational demands and the scarcity of annotated datasets limit their practicality for academic researchers. In this work, we introduce VideoLLaMB, a novel and efficient framework for long video understanding that leverages recurrent memory bridges and temporal memory tokens to enable seamless encoding of entire video sequences with preserved semantic continuity. Central to our approach is a SceneTiling algorithm that segments videos into coherent semantic units, facilitating robust understanding across tasks without requiring additional training. VideoLLaMB achieves state-of-the-art performance, surpassing existing models by 4.24.24.2 points on four VideoQA benchmarks and by 2.062.062.06 points on egocentric planning tasks. Notably, it maintains strong performance under extreme video length scaling (up to 8×8$\times$8 ×) and excels at fine-grained frame retrieval on our proposed Needle in a Video Haystack (NIAVH) benchmark. With linear GPU memory scaling, VideoLLaMB processes up to 320 frames using a single Nvidia A100 GPU, despite being trained on only 16 frames—offering an unprecedented balance of accuracy, scalability, and cost-effectiveness. This makes it highly accessible and practical for the academic community.
@@ 1 Introduction
Recent advancements in large-scale video language models, exemplified by GPT-4o and Project Astra have captivated global attention due to their potential for sophisticated interaction with real-world environments [23, 81]. These models are particularly noteworthy for their capacity to comprehend streaming video [65], which can be conceptualized as video with an unlimited context length. This capability necessitates both the observation of the current state and the ability to leverage long-term memory. Despite their promise, the training of such large-scale video-language (VidL) foundational models remains impractical for academic researchers. This impracticality arises from the substantial computational resources required by the complex, high-dimensional nature of long streaming video data, in addition to the scarcity of well-annotated, publicly available video-language datasets. These factors present significant challenges to scaling video-language models to the extent observed with large language models (LLMs).
@@
To circumvent these challenges, the community has witnessed a growing interest in developing computationally efficient multimodal large language models (MLLMs). Traditional methods resort to video compression strategies, such as sampling [76, 34], aggregation [70], semantic consolidation [54], and resampling [39, 20], in order to temporally reduce the length of the video. Yet, these methods often lead to the loss of critical visual cues, undermining the model’s ability to capture essential cues. Other approaches [47, 61] segment videos into shorter clips to mitigate the computational load of processing long videos. However, segmentation can disrupt the semantic flow of content, complicating the encoding process and potentially impacting the general understanding of the video narrative. Lastly, prevalent video understanding benchmarks, primarily based on linguistic question-answering pairs, exhibit static [26] and/or language biases [50, 82]. These biases favor models that rely more on static imagery or textual elements, respectively, and fail to provide a comprehensive assessment of a model’s capability on extended video sequences.
@@
Figure 1: Overview of VideoLLaMB. We first extract the video features using an off-the-shelf vision encoder, then apply SceneTiling to segment the video into semantic segments ($\S$2.1). Next, we use recurrent memory on these semantic segments to store video information within memory tokens ($\S$2.2). We further employ a retrieval mechanism to update the memory tokens and address long-dependency issues($\S$2.3). Finally, we project the memory-token-augmented features from the current video segment into the LLM.
@@
To address these multifaceted limitations, we introduce VideoLLaMB, an innovative framework that learns temporal Memory tokens within Bridge layers that recursively encode the entire video content, ensuring that no visual information is discarded deliberately (Figure 1; $\S$2). Specifically, we devise Memory Bridge Layers, equipped with recurrent memory tokens, that function without altering the architecture of the visual encoder and LLM. Furthermore, to mitigate the backpropagation through time (BPTT) issue, we maintain long-term dependencies by preserving recurrent memory tokens in a memory cache, which is periodically refreshed through a retrieval process. To compensate for the limitations of the sliding window technique, we propose the SceneTiling algorithm that divides the video into relatively independent sequences of semantic segments. This reduces the dimensions within each semantic unit without sacrificing semantic details. By constructing our recurrent memory with a retrieval mechanism based on these semantic segments, our method strikes a balance between effective and efficient comprehension of the current state and long-term memory retention.
@@
In $\S$3, we highlight the empirical advantages of VideoLLaMB in comparison with prior arts as:
  • Long-Term Memory Reservation. We demonstrate the effectiveness of VideoLLaMB in comprehensive long video understanding and enhanced frame retrieval through rigorous testing on established benchmarks. Using VideoMME [16], EgoSchema [41] and NexTQA [67], VideoLLaMB shows an average improvement of 4.24.24.2 accuracy points over PLLaVA [70], despite utilizing the same initialization and training video dataset. Notably, VideoLLaMB maintains robust performance even when video lengths extend to eight times their original duration. To further evaluate frame retrieval capabilities, we introduce the multimodal Needle in a Video Haystack (NIAVH) test, where VideoLLaMB consistently identifies the correct image within videos ranging from 1 to 320 seconds, surpassing other methods as video length increases.
@@
  • Real-time egocentric planning. To evaluate our model’s performance in video planning tasks, we used the dataset EgoPlan [9]. Our method achieves the best performance among all 7B video-language models, showing an improvement of 2.062.062.06 accuracy points over PLLaVA.
  • Training-free streaming captioning. By employing the SceneTiling algorithm, our method can automatically predict the end of a caption in streaming video without relying on special tokens during the training phase.
@@ 2 VideoLLaMB
VideoLLaMB is an extensible framework designed to enhance long video understanding, composed of three key modules: semantic-based segmenter ($\S$2.1), recurrent memory layer ($\S$2.2), and memory retriever ($\S$2.3). Each of these components will be detailed in the subsequent sections. Figure 1 depicts the overall framework.
@@ 2.1 SceneTiling: Segmentation with Semantics
Semantic segmentation along temporal sequence has long been recognized as an important task because it preserves the non-linear structure of context and greatly aids in compressing extensive context [48, 8, 43, 22, 60]. To address the disruption of semantic flow (see $\S$1), we introduce SceneTiling, a model-free scene segmentation algorithm inspired by TextTiling [21]. SceneTiling divides the entire video sequence into segments that are semantically distinct, ensuring intra-segment coherence.
@@
Aside from temporal semantic segmentation, SceneTiling enables streaming video captioning without requiring training with special tokens [7, 83, 15].
@@ 2.2 Recurrent Memory Bridge Layers
Traditional recurrent memory-based Transformers [4, 5, 25] incur significant computational costs when scaled up, i.e., $\mathcal{O}(L^{2}/K)$, where $L$ is the context length and $K$ is the number of segments, primarily due to its recurrent mechanism over the whole language model. More recently, some works empirically identify that linear projection best withstands visual information within MLLMs [35, 36, 79], albeit with high space complexity, whereas the resampler has strong compressing ability on semantic information [28], though it tends to miss detailed information [69].
In this work, we devised a novel Recurrent Memory Bridge Layer, implemented as Transformer blocks, that integrates recurrent memory tokens within bridge layers to enhance the linear layer’s memorization ability. Formally, for each video segment $s_{i}$, we prepend a fixed number of memory tokens, denoted as $[m_{i};s_{i}]$, where $m_{i}$ represents the memory tokens. Subsequently, we apply standard self-attention to this sequence, yielding $[m_{i+1};o_{i}]={\rm BridgeLayer}([m_{i};s_{i}])$. Here, $m_{i+1}$ is the updated memory token, and $o_{i}$ is the visual representation from the bridge layers. This process is carried out recursively, traversing the semantic video segments while updating the memory tokens. After a total of $k$ steps, this output represents the condensed visual summary of the video sequence and will serve as the input for the LLM. As such, the Memory Bridge can compress past video into memory while preserving current video scenes through projection without losing detailed information.
@@ 2.3 Memory Cache with Retrieval
One of the primary challenges associated with recurrent memory bridge layers is the potential for gradient vanishing, which can impede the model’s ability to learn long-range dependencies [29]. To mitigate this issue, we propose the incorporation of a memory cache with a retrieval strategy designed to preserve previous states of memory.
@@
At each timestep $i$, the system stores all previous memory tokens in a memory cache, denoted as $M_{i}=[m_{1},\ldots,m_{i}]$. We employ a self-retrieval mechanism to update the current memory token $m_{i}$. Specifically, we treat $m_{i}$ as a query and the concatenated memory cache $M_{i}$ as key and value. The model performs a standard multi-head cross-attention operation to integrate information from previous timesteps into the current memory state, yielding the updated memory token $m_{i+1}=\text{Softmax}\left({W_{i}^{Q}m_{i}(W_{i}^{K}M_{i})^{\top}}/{\sqrt{d_{k}}}\right)W_{i}^{V}M_{i}$, where $W_{i}^{Q},W_{i}^{K},W_{i}^{V}$ are weight matrices for query, key and value, respectively.
@@ 3.1 Setup
We utilize Vicuna-7B-v1.5 as the LLM and ViT-L/14 as the visual backbone following Video-LLaVA [34]. Each frame is resized and cropped to a dimension of 224×224. The Memory Bridge Layers are based on a single-layer Transformer. Our model is trained and evaluated with 16 frames and 4 segments, following the same video data protocol as PLLaVA [70].
@@ 3.2 Long-form Video Understanding
We conduct a comparative analysis of two types of models: retrieval-based methods and generative video-language models, as elaborated in Section 4. To ensure fairness in our comparisons, we primarily focus on state-of-the-art models such as LLaVA-NeXT-Video-DPO [79] and PLLaVA [70], which utilize the same base model and video datasets as our approach. Other SoTA models, including MovieChat [54], MA-LMM [20], VideoStreaming [47], and Video-xl [52], are not consistently included in all benchmarks due to variations in training data, model configurations, and benchmark settings. Nevertheless, we evaluate the key compression settings of these baselines using the same data and model configurations as ours, as detailed in Sections 3.6 and 3.8.
@@
Video-LLaVA 2024 Vicuna-7B 8 40.2
LLaVA-NeXT-Video-DPO 2024d Vicuna-7B 32 41.6
@@
Table 1: Results on Subset of EgoSchema under zero-shot setting. : the model has been fine-tuned using the training data from EgoSchema. $(n)$: $n$ frames are used in training.
@@
LLaMA-VID 2024d 53.8 60.0 73.0 59.5
Video-LLaVA 2024 56.9 61.0 75.0 61.3
LLaVA-NeXT-Video-DPO 2024d 55.6 61.0 73.9 61.3
@@
Method
Vision
Encoder
LLM
Size
AS AP AA FA UA OE OI OS MD AL ST AC MC MA SC FP CO EN ER CI Avg.
@@
LLaVA 2023 ViT-L 7B 28.0 39.5 63.0 30.5 39.0 53.0 41.0 41.5 23.0 20.5 45.0 34.0 20.5 38.5 47.0 25.0 36.0 27.0 26.5 42.0 36.0
Video-LLaMA 2023 CLIP-G 7B 27.5 25.5 51.0 29.0 39.0 48.0 40.5 38.0 22.5 22.5 43.0 34.0 22.5 32.5 45.5 32.5 40.0 30.0 21.0 37.0 34.1
LLaMA-Adapter 2024c ViT-B 7B 23.0 28.0 51.0 30.0 33.0 53.5 32.5 33.5 25.5 21.5 30.5 29.0 22.5 41.5 39.5 25.0 31.5 22.5 28.0 32.0 31.7
Video-ChatGPT 2024 ViT-L 7B 23.5 26.0 62.0 22.5 26.5 54.0 28.0 40.0 23.0 20.0 31.0 30.5 25.5 39.5 48.5 29.0 33.0 29.5 26.0 35.5 32.7
VideoChat 2023 CLIP-G 7B 33.5 26.5 56.0 33.5 40.5 53.0 40.5 30.0 25.5 27.0 48.5 35.0 20.5 42.5 46.0 26.5 41.0 23.5 23.5 36.0 35.5
VideoChat2β 2024c UMT-L 7B 66.0 47.5 83.5 49.5 60.0 58.0 71.5 42.5 23.0 23.0 88.5 39.0 42.0 58.5 44.0 49.0 36.5 35.0 40.5 65.5 51.1
@@
Table 3: Results on MVBench [32] multi-choice question answering. We highlight top-3 results among all 7B models of each category in purple. $\alpha$: training data from Xu et al. [70]. $\beta$: training with data from Li et al. [32].
Model Data Short Medium Long All. Comp.
LLaVA-NeXT-Vicuna 2024d LLaVA-NeXT 35.26 37.44 32.88 35.44 1.00
VideoChat2 2023 VideoChat2 - - - 33.7 0.49
PLLaVA 2024b PLLaVA 46.44 38.00 33.22 38.22 0.25
VideoLLaMB $\alpha$ (Ours) PLLaVA 46.11 38.44 34.22 39.59 0.06
VideoLLaMB $\beta$ (Ours) VideoChat2 49.22 39.11 35.89 41.41 0.06
Table 4: Results on VideoMME. We list models that use the same training data and Vicuna-7B backbones for fair comparison, for existing SoTA models like Video-XL [52], VideoStreaming [47] are trained on self-constructed dataset. Comp: Compression.
@@
EgoSchema [41] consists of egocentric videos, each averaging 180 seconds in length. This video QA dataset focuses on aspects such as understanding, reasoning, and long-term memory. In our experiment, we follow the precedent set by previous studies and use the public subset for evaluation. The results are presented in Table 2. Overall, our method significantly outperforms current generative video language models trained on similar data, demonstrating robust performance compared to other approaches and confirming its efficacy. Specifically, we compare our method with PLLaVA [70], which shares the same training data, LLM backbones, and input number of frames. Our method shows significant improvements over PLLaVA, indicating its superiority in understanding long egocentric videos. While our method does not yet match the performance of fine-tuned retrieval-based methods, we plan to apply our approach to larger language models to bridge this performance gap.
@@
The model is trained on 16-frame sequences, divided into 4 segments. However, in real-world scenarios, videos can be significantly longer than this training configuration. To demonstrate VideoLLaMB’s ability to extrapolate to longer videos, we conducted experiments on EgoSchema under two conditions: 1) dynamic segments, which adaptively control the number of segments based on the SceneTiling threshold, and 2) static segments, fixed at 4 segments. Results in Figure 2 reveal that dynamic segments are more effective than static segments, especially for shorter videos, indicating that our method can effectively maintain an appropriate number of segments. However, as video length increases, the performance of dynamic segments declines, notably at the 32-frame mark, where both strategies use four segments. Beyond this point, increasing the number of segments results in diminishing returns, likely due to the domain gap from training on shorter videos. To address this issue, we plan to fine-tune our models on longer videos for more substantial improvements. Overall, compared to PLLaVA, our method maintains consistent performance as the input length increases. In summary, our approach effectively extracts key information from videos, outperforming the simple pooling strategies used for memory consolidation in existing methods.
Results on NExTQA
@@
NExTQA [67], featuring daily-life videos that average 45 seconds in length, is designed to test a variety of question types, specifically temporal, causal, and descriptive questions. We applied our method to NExTQA to evaluate its temporal grounding ability. To maintain consistency with established benchmarks, we used the validation set for evaluation. In Table 2, we present the comprehensive results of our analysis. For a fair comparison, our primary benchmark is against PLLaVA, which includes instruction data from the NExTQA training set. Our method surpasses PLLaVA by 2.92.92.9 points. Notably, our approach demonstrates a significant enhancement in the temporal setting, achieving a 4.64.64.6 point improvement over PLLaVA. These results indicate that our scene-segment aware method effectively improves the model’s temporal grounding ability by compressing abundant information within scenes that share high semantic similarity.
Results on Long Context VideoQA
We evaluate our methods on the long video benchmark, VideoMME [16], which contain videos ranging from 11 seconds to 1 hour in length. The results are illustrated in Table 4. We compare methods using the same training data and LLM backbones. Compared to PLLaVA-B [70] and VideoChat-2 [30], our method shows improvements on both benchmarks. Notably, our method demonstrates consistent improvements on long videos in Video-MME with much less compression rate.
3.3 Comprehensive Video Understanding